OpenStudy (anonymous):

I just want to make sure, is radical -40 = 2i radical 10?

OpenStudy (danjs):

\[\sqrt{-40} = \sqrt{-1 * 40}\]

OpenStudy (danjs):

\[\sqrt{-1}\sqrt{4 * 10} = i*\sqrt{4}\sqrt{10} = 2i \sqrt{10}\]

OpenStudy (danjs):

yep you are righ

OpenStudy (mathstudent55):

Answer to your question: Yes, you are correct.

OpenStudy (anonymous):

thank you